For the reactants [NH3CH2COOH]+ and CH3CH2O- what is the deltaHrxn value? This is an acid base reaction that should be analyzed with pKa values to find deltaHrxn.
Answer #1

[NH3CH2COOH]+ is the protonated form of glycine (NH2CH2COOH)

The pKa value of glycine = 9.6 (for NH3+ group)

CH3CH2O- is the deprotonated form of ethanol (CH3CH2OH)

The pKa value of ethanol = 15.9 (for OH group)

The reaction between [NH3CH2COOH]+ and CH3CH2O- can be written as follows.

[NH3CH2COOH]+ + CH3CH2O- NH2CH2COOH + CH3CH2OH

The equilibrium shifts to the right side, i.e. forward reaction is favored. This is because of (pKa)protonated glycine < (pKa)ethanol,i.e. (Ka)protonated glycine > (Ka)ethanol

Since the reaction moves to the forward direction, the value of Hrxn is negative.

Here, pKeq = 9.6 - 15.9 = -6.3

i.e. -Log(Keq) = -6.3

i.e. The equilibrium constant (Keq) = 106.3 ~ 2*106

Now, The change in free energy of the reaction (Grxn) = -RT ln(Keq)

Where 'R' is the universal gas constant = 8.3145 J/mol.K

'T' is the absolute temperature = 25 oC (at STP) = (25+273.15) K = 298.15 K

i.e. Grxn = -8.3145 J/mol.K * 298.15 K * ln(2*106)

i.e. Grxn ~ -36 kJ/mol